OpenStudy (anonymous):

Solve for the variable. a-(-98)=-2 a=?

OpenStudy (ash2326):

We have \[\large a-(-98)=-2\] The - before the bracket will change the sign of the term inside. We already have a -98 there so it'll become 98 \[\large a+98=-2\] Now subtract both sides by 98 \[\large a+98-98=-2-98\] \[\large a+\cancel {98}-\cancel {98}=-2-98\] We have now \[\large a+0=-100\] or \[\huge a=-100\]

OpenStudy (anonymous):

Thanks:)

OpenStudy (ash2326):

You're welcome:D